OpenStudy (anonymous):

If it requires 7.0 J of work to stretch a particular spring by 2.3 cm from its equilibrium length, how much more work will be required to stretch it an additional 3.9 cm?

OpenStudy (anonymous):

w=f.x, so 7=f.2.3 , f= 7/2.3 => f= 3.04 energy of a spring is (1/2) k x^2 so k is spring const k= f/ x=> 3.04/2.3 = 1.32 extra energy will be (1/2)x1.32x (3.9)^2 i.e; 10 j i didn't consider units so plz re do this by considering the units